A 5.00g mixture contains both lithium fluoride, LiF, and potassium fluoride,KF. If the mixture contains 3.10g of fluorine, what is the mass of KF in the mixture
1
Expert's answer
2020-04-02T08:57:13-0400

Let the LiF be x then KF will be (5-x)

molar mass of LiF = 26

molar mas of KF = 58

moles of LiF = "\frac{x}{26}"

moles of KF ="\frac{5-x}{58}"


amount of total fluorine present = amount of fluorine present in (LiF +KF)

amount of fluorine present in LiF = "\frac{19}{26}\times\frac{x}{26}"


amount of fluorine present in KF ="\frac{19}{58}\times \frac{5-x}{58}"

total fluorine is equal to 3.10 grams

so lets substitute and get the value of x

"\frac{19}{26}\times\frac{x}{26}+\frac{19}{58}\times \frac{5-x}{58}=3.1"


"19x\times58+19\times26(5-x)=3.1\times58\times26\\"

"1102x+2470-494x=4674.8\\608x=2204.8\\x=3.626\ grams"


mass of KF="5-x=5-3.626=1.373 grams"